# From loop quantum gravity to cosmology: the 2-vertex model

### v2

@@ IV.3 Geometric interpretation
$\xi_{i}^{A}=i\ln\frac{(e^{-i\varphi/2}\bar{w}_{i}^{A})w_{i}^{A}}{(e^{i\varphi/2}w_{i}^{A})\bar{w}_{i}^{A}}=\varphi\mathrel{\mathop{\ordinarycolon}}=\xi.$ (34)
@@
In summary, we have found that the reduced sector (under the global U($N$) symmetry) of the 2-vertex model depends solely on two canonically conjugate variables which are homogeneous and isotropic, i.e. independent of the edges and vertices. The first one is the total area $A$ of the polyhedron, directly related to its volume through (32). Therefore, its evolution can be interpreted as expansion. The second phase space variable is the twist angle $\varphi=\xi$ which, as briefly discussed in the next paragraph, is related to the extrinsic curvature. Thus, our model possesses a unique curvature that is independent of edges and a single polyhedron-area independent of vertices. Considering the results and conclusions of previous works [14, 15], the fact that our dynamical variables admit an interpretation of area (one-to-one related with the volume by Eq. (32)) and extrinsic curvature strengthens the cosmological interpretation of this model.
According to the proposal of Freidel and Speziale [24], the discrete twist angle $\xi_{i}$ associated with the edge $i$ can be related to the continuous extrinsic curvature by $\xi_{i}^{2}=-\ell^{2}\det(\hat{\ell}_{i}^{a}K_{a}^{I}\sigma_{I})$, where $\ell$ is a small length scale, ^$\hat{\ell}_{i}^{a}$ is a unitary vector along the edge $i$, and $K_{a}$ is the $\mathfrak{su}(2)$-valued extrinsic-curvature one-form. In the U$(N)$-symmetric 2-vertex model, the twist angle does not depend on the edge $i$ as we have seen and hence the edge information disappears from the relation between twist angle and the extrinsic curvature proposed in [24] pointing towards an homogeneity and isotropy, which we pin down in the following section.
V LQC from the U$(N)$-invariant 2-vertex model
In this section we show that the U$(N)$-invariant 2-vertex model is indeed a reduction of the full LQG theory which provides LQC effective dynamics.
@@ V.1 The constraint equation
Let us start by noting that the 2-vertex Hamiltonian (12) can be modified as long as we respect the SU(2) symmetry and preserve the commutation with the matching constraint. Thus, we consider the following generalization of the U($N$)-reduced Hamiltonian constraint
@@
V.2 FLRW cosmology as the small-twist regime of the 2-vertex model
@@
Note that in this limit the Barbero-Immirzi parameter disappears as expected. From this Hamiltonian (or equivalently from (38)) we obtain $\dot{a}=-\mathcal{N}\gamma{\pi_{a}}/(2a)$, which replaced back into the constraint, yields
@@
where $\mathcal{N}(t)$ is the lapse function, $a(t)$ is the scale factor, $\mathop{}\!\mathrm{d}\Sigma_{k}^{2}={\mathop{}\!\mathrm{d}r^{2}}/({1-kr^{2}})+r^{2}\mathop{}\!\mathrm{d}\Omega_{2}^{2}$, $k$ is the constant that characterizes the curvature of the spatial slices, and $\mathop{}\!\mathrm{d}\Omega_{2}^{2}$ is the metric of the unit 2-sphere. The general-relativistic dynamics for this FLRW universe filled with a perfect fluid with density $\rho$ and pressure $p$ is described by the Friedmann’s equation
@@
In order to compare both dynamical equations (42) and (44) on equal footing it is necessary to introduce a fiducial volume $\mathcal{V}_{\Sigma}$ with respect to the line element $\mathop{}\!\mathrm{d}\Sigma_{k}^{2}$ (see footnote 3). This factor naturally appears multiplying both $\rho$ and $1/G$ when considering the Hamiltonian for FLRW dynamics. In this comparison we then see that the small-twist regime of the U$(N)$-invariant 2-vertex model describes a FLRW cosmology. The cosmological parameters and energy density are determined by the parameters $\gamma$, $\kappa$, and the function $f$ that define the 2-vertex model: The gravitational constant is given by $G=3\gamma\mathcal{V}_{\Sigma}/(8\pi)$, the spatial curvature by $k=\kappa$, and the matter content is a barotropic fluid with density $\rho=f/\mathcal{V}_{\Sigma}$ and pressure determined by the continuity equation $p=-(a^{3}f)^{\prime}/(3\mathcal{V}_{\Sigma}a^{2})$, the prime denoting derivative with respect to $a$.
As a simple example, the 2-vertex model with
@@
Therefore, Freidel and Speziale’s interpretation of the twist angle in terms of the extrinsic curvature [24] finds here an explicit realization which allows in addition to interpret $\gamma$ (and hence $G/(\beta^{2}\mathcal{V}_{\Sigma})$) as the inverse of the fundamental length scale in the 2-vertex model
@@ V.3 LQC effective dynamics
In the previous section we have shown that the 2-vertex dynamics describes an FLRW spacetime filled with perfect barotropic fluids in the lowest order in the twist angle, that is, in the low-curvature regime. This has allowed us to write the gravitational constant, the spatial curvature, and the matter density in terms of the parameters $\gamma$ and $\kappa$ and the function $f$ that characterize the 2-vertex model.
In the old dynamics of LQC for a universe with a free scalar field and vanishing spatial curvature, the semiclassical states (i.e. Gaussian coherent states sharply peaked at the classical FLRW trajectories in the low-curvature regime) follow the effective Friedmann’s equation [35]
@@
Comparing Eq. (39) with $\kappa=0$ (as corresponds to vanishing spatial curvature) and (51) we observe that the U($N$)-symmetric 2-vertex model describes the old effective dynamics of LQC with an area gap $\Delta=1/(2\gamma)^{2}$ and a minimum fiducial length $\mu_{0}=\sqrt{\Delta}=1/(2\gamma)$.
@@ V.4 Towards LQC improved dynamics
The U($N$)-symmetric 2-vertex model describes the old effective dynamics of LQC. Nevertheless, we can try to implement the improved dynamics [36] in a similar way to that of LQC by suitably modifying the model.
To do so, let us go back to the relation (23) imposed by the U($N$) symmetry on the spinors. Note that, in the U($N$) sector, the Poisson structure (3) for the spinors determines the Poisson bracket between the twist angle $\varphi$ and the total area $A$, so that they are canonically conjugate variables. To introduce the improved dynamics, we need to change this structure so that $\varphi$ and $A$ are no longer canonically conjugate. One way of doing this is to change the Poisson bracket to $\{\varphi,\sqrt{A}\}=1$, while keeping the Hamiltonian (35). Another equivalent option, which is the one we will use in the following, is to introduce a new variable $\phi=\varphi/\sqrt{A}$ canonically conjugate to $A$, so that $\{\phi,A\}=1$. Then, the U($N$) relation (23) now becomes
$|w_{i}\rangle=e^{-i\phi\sqrt{A}}|\bar{z}_{i}\rangle.$ (52)
The spinor Poisson structure that leads to this modification is no longer (3), so this modified model is not a direct consequence of the 2-vertex model but of a suitably modified one, whose structure is beyond the scope of this paper. The Hamiltonian in terms of these new canonical variables $\phi$ and $A$ is
$\breve{H}=2\mathcal{M}A^{2}\gamma\big{[}-1+\cos(\sqrt{A}\phi)+g(A)\big{]},$ (53)
@@
If we now follow steps analogous to those of Sec. V.3, we obtain
$\frac{1}{N^{2}}\frac{\dot{a}^{2}}{a^{2}}=\gamma f(a)\Big{[}1-\frac{\beta^{3}}{4\gamma}f(a)\Big{]}.$ (54)
On the other hand, the improved dynamics formalism introduces in the holonomies a non-constant minimum fiducial length $\bar{\mu}=\sqrt{\Delta}/a$. The resulting effective LQC equation is [37]
$\frac{1}{\mathcal{N}^{2}}\frac{\dot{a}^{2}}{a^{2}}=\frac{8\pi G}{3}\rho\Big{(}1-\frac{\rho}{\rho_{\star}}\Big{)},$ (55)
where now the maximum energy density (reached at the bounce) is precisely $\rho_{\star}$ and, thus, independent of the initial energy density $\rho_{0}$. We see that the modified U($N$)-symmetric 2-vertex dynamics (54) provides the effective improved dynamics (55) with energy density $\rho=f/\mathcal{V}_{\Sigma}$, gravitational constant $G=3\gamma\mathcal{V}_{\Sigma}/(8\pi)$ and area gap $\Delta=\beta/(4\gamma^{2})$. Consequently, the $\bar{\mu}$ of the LQC improved dynamics can be written in terms of the modified 2-vertex parameters as $\bar{\mu}=\sqrt{\Delta}/a=\sqrt{\beta}/(2\gamma a)$.
As we have already mentioned, this result is a consequence of an ad hoc modification of the 2-vertex model (very much in the same way as the improved dynamics is introduced in LQC). Nevertheless, this modification is useful to further understand the underlying physics behind the imposition of the improved dynamics in LQC and its relation with full LQG.
@@ VI Conclusions
The 2-vertex graph has proved to be a useful model within LQG in order to implement dynamics and to study the emergence of an interesting cosmological behavior, at least in a symmetry-reduced sector [14, 15, 16, 17, 18]. Therefore, this model provides an excellent arena to explore the main open problems of the theory. On the other hand, the spinorial formalism introduced in [19] gives us a convenient mathematical formulation of the kinematic Hilbert space of LQG.
@@
We have made use of the spinorial formalism applied to the 2-vertex model in order to analytically study the evolution of the classical volume. This formalism provides normal vectors to the faces of the closed polyhedra (the conditions for the Minkowski theorem are constraints) associated with each of the vertices. Nevertheless, there is no known analytical formula to compute the volume of a general polyhedron (with $N>4$ faces) in terms of the normal vectors that generate it. Instead, we have used geometric quadrupoles [27] in order to construct an approximation of the volume that reproduces the qualitative behavior of the volume [17]. We have calculated the equation of motion for this approximation, finding that degenerate polyhedra—with all the faces parallel to each other—have null volume forever, i.e., zero-volume singularities are decoupled from the rest of configurations.
@@
The $\text{U}(N)$-symmetric sector of the 2-vertex model contains just one degree of freedom corresponding to the twist angle and the total area of the polyhedra, which are canonically conjugate. We have found that the approximation to the volume in terms of quadrupoles is proportional to the exact volume for the case of the tetrahedron ($N=4$). It is left for future work to check whether this behavior is true for any other number of faces. Comparing the evolution of the quadrupole approximation to the volume and the total area in the $\text{U}(N)$-symmetric sector, we have found that they precisely follow the evolution one expects from a homogeneous and isotropic evolution. This relation is yet another argument for the use of the quadrupole approximation of the volume and for the interpretation of the 2-vertex model as a cosmological model (at least in the $\text{U}(N)$-reduced sector).
@@
We have shown that a generalization of the 2-vertex Hamiltonian constraint (that preserves the symmetries of the model) consisting of the addition of a function of the area, effectively takes into account matter content in the form of arbitrary barotropic perfect fluids. In particular, we have shown that there is a direct relation between the equation of state and the specific function of the area introduced in the 2-vertex Hamiltonian.
@@
Finally, we have proved that the 2-vertex model is indeed an LQG truncation that effectively describes LQC with arbitrary barotropic perfect fluids. The LQC flavor that derives directly from the 2-vertex model is the old dynamics. However, a suitable modification of the Poisson bracket structure so that the twist angle is canonically conjugate to the square root of the area (and not the area itself) leads to the LQC improved dynamics.
@@ Acknowledgements.
We would like to thank Etera R. Livine and Mercedes Martín-Benito for very enlightening conversations. ÁC acknowledges financial support from IPARCOS Ayudas de Máster IPARCOS-UCM/2022 and from the Spanish Government fellowship FPU22/03222. This work was supported by the Basque Government Grant IT1628-22, and by the Grant PID2021-123226NB-I00 (funded by MCIN/AEI/10.13039/501100011033 and by “ERDF A way of making Europe”). This work was also supported by the Grant PID2020-118159GB-C44 (funded by MCIN/AEI/10.13039/501100011033).

### v3

@@ IV.3 Geometric interpretation
$\xi_{i}^{A}=i\ln\frac{(e^{-i\varphi/2}\bar{w}_{i}^{A})w_{i}^{A}}{(e^{i\varphi/2}w_{i}^{A})\bar{w}_{i}^{A}}=\varphi:=\xi.$ (34)
@@
In summary, we have found that the reduced sector (under the global U($N$) symmetry) of the two-vertex model depends solely on two canonically conjugate variables which are homogeneous and isotropic, i.e. independent of the edges and vertices. The first one is the total area $A$ of the polyhedron, directly related to its volume through (32). Therefore, its evolution can be interpreted as expansion. The second phase space variable is the twist angle $\varphi=\xi$ which, as briefly discussed in the next paragraph, is related to the extrinsic curvature. Thus, our model possesses a unique curvature that is independent of edges and a single polyhedron-area independent of vertices. Considering the results and conclusions of previous works [14, 15], the fact that our dynamical variables admit an interpretation of area (one-to-one related with the volume by Eq. (32)) and extrinsic curvature strengthens the cosmological interpretation of this model.
According to the proposal of Freidel and Speziale [24], the discrete twist angle $\xi_{i}$ associated with the edge $i$ can be related to the continuous extrinsic curvature by $\xi_{i}^{2}=-\ell^{2}\det(\hat{\ell}_{i}^{a}K_{a}^{I}\sigma_{I})$, where $\ell$ is a small length scale, ^$\hat{\ell}_{i}^{a}$ is a unitary vector along the edge $i$, and $K_{a}$ is the $\mathfrak{su}(2)$-valued extrinsic-curvature one-form. In the U$(N)$-symmetric two-vertex model, the twist angle does not depend on the edge $i$ as we have seen and hence the edge information disappears from the relation between twist angle and the extrinsic curvature proposed in [24] pointing towards an homogeneity and isotropy, which we pin down in the following section.
V LQC from the U$(N)$-invariant two-vertex model
In this section we show that the U$(N)$-invariant two-vertex model is indeed a reduction of the full LQG theory which provides LQC effective dynamics.
@@ V.1 The constraint equation
Let us start by noting that the two-vertex Hamiltonian (12) can be modified as long as we respect the SU(2) symmetry and preserve the commutation with the matching constraint. Thus, we consider the following generalization of the U($N$)-reduced Hamiltonian constraint
@@
V.2 FLRW cosmology as the small-twist regime of the two-vertex model
@@
Note that in this limit the Barbero-Immirzi parameter disappears as expected. From this Hamiltonian (or equivalently from (38)) we obtain $\dot{a}=-\mathcal{N}\gamma{\pi_{a}}/(2a)$, which, when replaced back into the constraint, yields
@@
where $\mathcal{N}(t)$ is the lapse function, $a(t)$ is the scale factor, $\mathop{}\!\mathrm{d}\Sigma_{k}^{2}={\mathop{}\!\mathrm{d}r^{2}}/({1-kr^{2}})+r^{2}\mathop{}\!\mathrm{d}\Omega_{2}^{2}$, $k$ is the constant that characterizes the curvature of the spatial slices, and $\mathop{}\!\mathrm{d}\Omega_{2}^{2}$ is the metric of the unit two-sphere. The general-relativistic dynamics for this FLRW universe filled with a perfect fluid with density $\rho$ and pressure $p$ is described by the Friedmann’s equation
@@
In order to compare both dynamical equations (42) and (44) on equal footing it is necessary to introduce a fiducial volume $\mathcal{V}_{\Sigma}$ with respect to the line element $\mathop{}\!\mathrm{d}\Sigma_{k}^{2}$ (see footnote 3). This factor naturally appears multiplying both $\rho$ and $1/G$ when considering the Hamiltonian for FLRW dynamics. In this comparison we then see that the small-twist regime of the U$(N)$-invariant two-vertex model describes a FLRW cosmology. The cosmological parameters and energy density are determined by the parameters $\gamma$, $\kappa$, and the function $f$ that define the two-vertex model: The gravitational constant is given by $G=3\gamma\mathcal{V}_{\Sigma}/(8\pi)$, the spatial curvature by $k=\kappa$, and the matter content is a barotropic fluid with density $\rho=f/\mathcal{V}_{\Sigma}$ and pressure determined by the continuity equation $p=-(a^{3}f)^{\prime}/(3\mathcal{V}_{\Sigma}a^{2})$, the prime denoting derivative with respect to $a$.
As a simple example, the two-vertex model with
@@
Therefore, Freidel and Speziale’s interpretation of the twist angle in terms of the extrinsic curvature [24] finds here an explicit realization which allows in addition to interpret $\gamma$ (and hence $G/(\beta^{2}\mathcal{V}_{\Sigma})$) as the inverse of the fundamental length scale in the two-vertex model
@@ V.3 LQC effective dynamics
In the previous section we have shown that the two-vertex dynamics describes an FLRW spacetime filled with perfect barotropic fluids in the lowest order in the twist angle, that is, in the low-curvature regime. This has allowed us to write the gravitational constant, the spatial curvature, and the matter density in terms of the parameters $\gamma$ and $\kappa$ and the function $f$ that characterize the two-vertex model.
In the old dynamics of LQC for a universe with a free scalar field and vanishing spatial curvature, the semiclassical states (i.e. Gaussian coherent states sharply peaked at the classical FLRW trajectories in the low-curvature regime) follow the effective Friedmann’s equation, as can be derived from [35],
@@
Comparing Eq. (39) with $\kappa=0$ (as corresponds to vanishing spatial curvature) and (51) we observe that the U($N$)-symmetric two-vertex model describes the old effective dynamics of LQC with an area gap $\Delta=1/(2\gamma)^{2}$ and a minimum fiducial length $\mu_{0}=\sqrt{\Delta}=1/(2\gamma)$.
@@ V.4 Towards LQC improved dynamics
The U($N$)-symmetric two-vertex model describes the old effective dynamics of LQC. Nevertheless, we can try to implement the improved dynamics [36] in a similar way to that of LQC by suitably modifying the model.
To do so, let us go back to the relation (23) imposed by the U($N$) symmetry on the spinors. Note that, in the U($N$) sector, the Poisson structure (3) for the spinors determines the Poisson bracket between the twist angle $\varphi$ and the total area $A$, so that they are canonically conjugate variables. To introduce the improved dynamics, we need to change this structure so that $\varphi$ and $A$ are no longer canonically conjugate. One way of doing this is to change the Poisson bracket to $\{\varphi,2\sqrt{A}\}=1$, while keeping the Hamiltonian (35). Another equivalent option, which is the one we will use in the following, is to introduce a new variable $\phi=\varphi/\sqrt{A}$ canonically conjugate to $A$, so that $\{\phi,A\}=1$. Then, the U($N$) relation (23) now becomes
$|w_{i}\rangle=e^{-i\phi\sqrt{A}/2}|\bar{z}_{i}\rangle.$ (52)
The spinor Poisson structure that leads to this modification is no longer (3), so this modified model is not a direct consequence of the two-vertex model but of a suitably modified one, whose structure is beyond the scope of this paper. The Hamiltonian in terms of these new canonical variables $\phi$ and $A$ is
$\breve{H}=2\mathcal{M}A^{2}\left[\gamma\left(-1+\cos\big{(}\sqrt{A}\phi\big{)}\right)+g(A)\right],$ (53)
@@
If we now follow steps analogous to those of Sec. V.3, then we obtain
$\frac{1}{N^{2}}\frac{\dot{a}^{2}}{a^{2}}=\gamma f(a)\Big{[}1-\frac{\beta^{3}}{4\gamma}f(a)\Big{]},$ (54)
where now $f(a)=2\beta^{-3}g(a^{2}/\beta)$.
On the other hand, the improved dynamics formalism introduces in the holonomies a nonconstant minimum fiducial length $\bar{\mu}=\sqrt{\Delta}/a$. The resulting effective LQC equation is [37]
$\frac{1}{\mathcal{N}^{2}}\frac{\dot{a}^{2}}{a^{2}}=\frac{8\pi G}{3}\rho\Big{(}1-\frac{\rho}{\rho_{\star}}\Big{)}.$ (55)
Therefore, now the maximum energy density (reached at the bounce) is precisely $\rho_{\star}$ and, thus, independent of the initial energy density $\rho_{0}$. We see that the modified U($N$)-symmetric two-vertex dynamics (54) provides the effective improved dynamics (55) with energy density $\rho=f/\mathcal{V}_{\Sigma}$, gravitational constant $G=3\gamma\mathcal{V}_{\Sigma}/(8\pi)$ and area gap $\Delta=\beta/(4\gamma^{2})$. Consequently, the $\bar{\mu}$ of the LQC improved dynamics can be written in terms of the modified two-vertex parameters as $\bar{\mu}=\sqrt{\Delta}/a=\sqrt{\beta}/(2\gamma a)$.
As we have already mentioned, this result is a consequence of an ad hoc modification of the two-vertex model (very much in the same way as the improved dynamics is introduced in LQC). Nevertheless, this modification is useful to further understand the underlying physics behind the imposition of the improved dynamics in LQC and its relation with full LQG.
@@ VI Conclusions
The two-vertex graph has proved to be a useful model within LQG in order to implement dynamics and to study the emergence of an interesting cosmological behavior, at least in a symmetry-reduced sector [14, 15, 16, 17, 18]. Therefore, this model provides an excellent arena to explore the main open problems of the theory. On the other hand, the spinorial formalism introduced in [19] gives us a convenient mathematical formulation of the kinematic Hilbert space of LQG.
@@
We have made use of the spinorial formalism applied to the two-vertex model in order to analytically study the evolution of the classical volume. This formalism provides normal vectors to the faces of the closed polyhedra (the conditions for the Minkowski theorem are constraints) associated with each of the vertices. Nevertheless, there is no known analytical formula to compute the volume of a general polyhedron (with $N>4$ faces) in terms of the normal vectors that generate it. Instead, we have used geometric quadrupoles [27] in order to construct an approximation of the volume that reproduces the qualitative behavior of the volume [17]. We have calculated the equation of motion for this approximation, finding that degenerate polyhedra—with all the faces parallel to each other—have null volume forever, i.e., zero-volume singularities are decoupled from the rest of configurations.
@@
The $\text{U}(N)$-symmetric sector of the two-vertex model contains just one degree of freedom corresponding to the twist angle and the total area of the polyhedra, which are canonically conjugate. We have found that the approximation to the volume in terms of quadrupoles is proportional to the exact volume for the case of the tetrahedron ($N=4$). It is left for future work to check whether this behavior is true for any other number of faces. Comparing the evolution of the quadrupole approximation to the volume and the total area in the $\text{U}(N)$-symmetric sector, we have found that they precisely follow the evolution one expects from a homogeneous and isotropic evolution. This relation is yet another argument for the use of the quadrupole approximation of the volume and for the interpretation of the two-vertex model as a cosmological model (at least in the $\text{U}(N)$-reduced sector).
@@
We have shown that a generalization of the two-vertex Hamiltonian constraint (that preserves the symmetries of the model) consisting of the addition of a function of the area, effectively takes into account matter content in the form of arbitrary barotropic perfect fluids. In particular, we have shown that there is a direct relation between the equation of state and the specific function of the area introduced in the two-vertex Hamiltonian.
@@
Finally, we have proved that the two-vertex model is indeed an LQG truncation that effectively describes LQC with arbitrary barotropic perfect fluids. The LQC flavor that derives directly from the two-vertex model is the old dynamics. However, a suitable modification of the Poisson bracket structure so that the twist angle is canonically conjugate to the square root of the area (and not the area itself) leads to the LQC improved dynamics.
@@ Acknowledgements.
We would like to thank Etera R. Livine and Mercedes Martín-Benito for very enlightening conversations. This work was partially supported by Grants No. PID2020–118159GB-C44, No. PID2021-123226NB-I00 and No. PID2022-139841NB-I00 (funded by MCIN/AEI/10.13039/501100011033 and by “ERDF A way of making Europe”), and the Basque Government Grant No. IT1628-22. Additionally, Á. C. acknowledges financial support from MIU (Ministerio de Universidades, Spain) fellowship FPU22/03222 and IPARCOS Ayudas de Máster IPARCOS-UCM/2022.